there is an available 75 liters of 55% solution of acid and water. what volume of water must be added to the solution to reduce the concentration to 18%?
Answer by stanbon(75887)

there is an available 75 liters of 55% solution of acid and water. what volume of water must be added to the solution to reduce the concentration to 18%?
----
Equation:
water + water = water
-----
0.45*75 + 1.00x = 0.82(75+x)
-------
45*75 + 100x = 82*75 + 82x
----
18x = 37
x = 2.056 liters (amt. of water needed)
===============
Cheers,
Stan H.